OpenStudy (anonymous):

solve exactly: 5^(x+4) = 1/125 x= ?

OpenStudy (anonymous):

1/125 is the same as 5^(-3)

OpenStudy (anonymous):

So x + 4 = -3

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

Do you see why 1/125 = 5^(-3)

OpenStudy (anonymous):

isnt that the same as 5^1/3

OpenStudy (anonymous):

No, fractional exponents gets into taking roots.

OpenStudy (anonymous):

A negative exponent means that you take 1/ So if you have x^(-n) that equals 1/x^n \[x^{-n} = \frac{1}{x^n}\]

OpenStudy (anonymous):

\[\frac{1}{125} = \frac{1}{5^3} = 5^{-3}\]

OpenStudy (anonymous):

that is right thank you

OpenStudy (anonymous):

You're welcome.